OpenStudy (anonymous):

use the logarithm properties to rewrite ln((cubedrt(c))/ab) in terms of r s and t where r=lna, b=lns, and t=lnc

OpenStudy (anonymous):

\[\ln(\frac{\sqrt[3]{c}}{ab})\] \[=\ln(\sqrt[3]{c})-\ln(ab)=\frac{1}{3}\ln(c)-\ln(a)-\ln(b)\] now substitute the numbers you know

OpenStudy (anonymous):

or i should say substitute the letter you are given. hope it is clear that \[\ln(\sqrt[3]{c})=\ln(c^{\frac{1}{3}})=\frac{1}{3}\ln(c)\]